OpenStudy (anonymous):

If h(x+2)=1/2 x , then h(-2) is

OpenStudy (raden):

just plug x=-4 to both sides what u get

OpenStudy (anonymous):

how you get -4

OpenStudy (anonymous):

how would i show my work for this problem

OpenStudy (raden):

h(x+2), if we plug x=-4 it would be h(-2), right ?

OpenStudy (anonymous):

yes than wat would i do next

OpenStudy (raden):

apply x=-4 to right side also, what u get ?

OpenStudy (anonymous):

-2 could you show me how to write it out solving it

OpenStudy (raden):

h(x+2)=1/2 x plug x= -4, h(-4+2) = 1/2 * (-4) h(-2) = -2 u are right

OpenStudy (anonymous):

is -2 the answer than

OpenStudy (raden):

sure...

OpenStudy (anonymous):

is that right or do i have to do another step

OpenStudy (raden):

well, we have another way to solve it.... h(x+2)=1/2 x let u=x+2, it means x=u-2 therefore, we get : h(u) = 1/2 * (u-2) now, if u we change again to x, gives h(x) = 1/2 * (x-2) now, plug x=-2 h(-2) = 1/2 * (-2-2) h(-2) = 1/2 * -4 = -2

OpenStudy (anonymous):

is -2 still the answer thou

OpenStudy (raden):

yes

OpenStudy (anonymous):

ok thanks can yu help with another problem as well

OpenStudy (raden):

sure..

OpenStudy (anonymous):

ok